OpenStudy (anonymous):

is 3/5 equal to or greater than 5/7

OpenStudy (jdoe0001):

well... let us multiply each fraction by a common value top and bottom, say \(\bf \cfrac{{\color{brown}{ 7}}\cdot 3}{{\color{brown}{ 7}}\cdot 5}\implies ?\qquad \cfrac{{\color{brown}{ 5}}\cdot 5}{{\color{brown}{ 5}}\cdot 7}\implies ?\)